OpenStudy (anonymous):

Find y' of y=log10 x (10 is the base)

OpenStudy (johnweldon1993):

\[\log_{10} x \] like that?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

\[\frac{ d }{ dx } \log_b (x) = \frac{ 1 }{ x \ln(b) }\] you have 10 as being the base...so in this equation b = 10

OpenStudy (anonymous):

Thank you !

OpenStudy (johnweldon1993):

No problem!

OpenStudy (dan815):

ur welcome

OpenStudy (jhannybean):

lmao.